OpenStudy (anonymous):

Faces: 25 Vertices: 17 Edges: ? 43 41 39 40

OpenStudy (anonymous):

Euler's formula states that V + F - E = 2

OpenStudy (anonymous):

so it would be like 17+25-E=2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

42-40=2

OpenStudy (anonymous):

Am I right?

Directrix (directrix):

Yes. V + F = E + 2 E = 40

OpenStudy (anonymous):

Vertices: 11 Edges: 34 Faces: ? (1 point) 25 28 26 24

OpenStudy (anonymous):

11+F+34=2 A.25 right?

OpenStudy (anonymous):

Edges: 36 Faces: 22 Vertices: ? (1 point) 19 15 16 17

OpenStudy (anonymous):

V+22-36=2 C.16 right

Directrix (directrix):

A.25 right? That is what I got.

OpenStudy (anonymous):

Mario’s company makes unusually shaped imitation gemstones. One gemstone had 12 faces and 10 vertices. How many edges did the gemstone have? (1 point) 23 edges 22 edges 25 edges 20 edges

OpenStudy (anonymous):

10+12-E=2 D.20

OpenStudy (anonymous):

Am i doing these right? @Directrix

Directrix (directrix):

D.20 I got this.

Directrix (directrix):

Any more?